### v2

# Vikhr: The Family of Open-Source Instruction-Tuned Large Language Models for Russian

Abstract
There has been a surge in the development of various Large Language Models (LLMs). However, text generation for languages other than English often faces significant challenges, including poor generation quality and the reduced computational performance due to the disproportionate representation of tokens in model’s vocabulary. In this work, we address these issues and introduce Vikhr, a new state-of-the-art open-source instruction-tuned LLM designed specifically for the Russian language. “Vikhr” refers to the name of the Mistral LLM series and means “strong gust of wind.” Unlike previous efforts for Russian that utilize computationally inexpensive LoRA adapters on top of English-oriented models, Vikhr features an adapted tokenizer vocabulary and undergoes the continued pre-training and instruction tuning of all weights. This approach not only enhances the model’s performance but also significantly improves its computational and contextual efficiency. The remarkable performance of Vikhr across various Russian-language benchmarks can also be attributed to our efforts in expanding instruction datasets and corpora for continued pre-training. Vikhr not only sets the new state of the art among open-source LLMs for Russian, but even outperforms some proprietary closed-source models on certain benchmarks. The model weights, instruction sets, and code are publicly available [^1] https://huggingface.co/Vikhrmodels.
1 Introduction
Instruction tuning has unlocked in Large Language Models (LLMs) vast zero-shot capabilities without the need of careful prompt engineering Ouyang et al. (2022). The most rapid research and development efforts are currently devoted to English LLMs. There has been a surge in English open-source models: Llama series Touvron et al. (2023a, b), Mistral series Jiang et al. (2023), Vicuna series Chiang et al. (2023), etc. This growth is driven by the abundance of raw training data in English and dedicated efforts to create comprehensive sets of instruction-output pairs. Despite the fact that LLMs oriented on English have some multilingual capabilities Zhao et al. (2024) due to small portions of texts in various languages leaked into their training datasets Touvron et al. (2023a), their overall performance in these languages remains relatively low. Although they can usually generate portions of coherent texts, these models struggle with reasoning in non-English languages, lack culture-specific knowledge, and are highly inefficient in terms of tokenization. This inefficiency arises due to the way bite-pair tokenization algorithms work: they split the infrequent words into multiple tokens. Since multilingual data typically represents a small portion of the training dataset, non-English words are often split in many pieces. This leads to more steps during prompt processing and text generation, shorter effective context windows, and ultimately lower quality Tikhomirov and Chernyshev (2023); Petrov et al. (2024). This disparity places non-English languages at a disadvantage.
There is a research direction focused on developing multilingual LLMs that work well for multiple popular languages: BLOOMz Muennighoff et al. (2023), mGPT Shliazhko et al. (2022), Bactrian-X Li et al. (2023), PALO Maaz et al. (2024), Aya101 from CohereAI Üstün et al. (2024), etc. These models are typically trained on rich multilingual datasets and are less skewed towards English. However, when aiming to perform well across multiple languages simultaneously, these models must still share their vocabulary and parameters. This often hinders their performance for each particular language in isolation, especially for the popular smaller model sizes, such as 7B and 13B.
The goal of maximizing the LLM performance for a specific language within a certain number of parameters has led researchers to develop bi-lingual LLMs. For example, Jais Sengupta et al. (2023) focus only on English and Arabic. The inclusion of English data in pre-training alongside Arabic data is motivated by the significantly larger volume of English data available. This helps LLMs substantially enhance skills such as logical and common sense reasoning, which are also applied when generating text in Arabic.
Russian is one of the high-resource languages and is typically represented in multilingual LLMs. Additionally, there are several proprietary closed-source LLMs, such as MTS AI, GigaChat, and YandexGPT, that meet or even surpass their English-oriented flagship competitors when it comes to text processing and generation in Russian. However, controllable research often requires white-box access to LLM logits and layer outputs, the ability to modify weights and a model architecture, and consistent answers for reproducibility, which is often impossible in closed-source LLMs due to their constant development and retirement. There are only a few open-source LLMs designed for Russian (such as Saiga Gusev (2023), ruGPT AI Forever (2022), ruadapt Tikhomirov and Chernyshev (2023)). Of these, only Saiga and ruadapt are instruction-tuned.
This work aims to build an efficient and effective open-source instruction-following LLM for Russian facilitating multilingual natural language processing research. Building even a small LLM that targets a particular language from scratch requires a lot of computational resources. Consequently, many researchers simply fine-tune LoRA adapters Hu et al. (2021) for English-oriented LLMs on some language-specific data. While this approach can improve model generation quality, it does not address computational inefficiency because the tokenizer and model vocabulary remain unchanged. In contrast, our approach not only fine-tunes a base LLM on Russian language data but also reconstructs its underlying tokenizer and vocabulary, alongside suggesting an improved method for continued pre-training. Additionally, we have significantly expanded the available Russian datasets for instruction tuning. The developed LLM achieves state-of-the-art results for the Russian language among other open-source counterparts across a wide range of benchmarks.
Contributions of the paper are the following:
  • We have constructed Vikhr – a state-of-the-art open-source instruction-following LLM oriented on the Russian language. In addition to its high generation quality, Vikhr features an efficient tokenizer that enables rapid text generation and good context utilization.
  • We have developed a pipeline for adapting English-oriented LLMs to the Russian language. The pipeline implements vocabulary adaptation, continued pre-training with regularization to prevent “catastrophic forgetting”, and instruction tuning.
  • We have expanded the datasets for continued pre-training of Russian language models and previously available instruction datasets.
  • We conducted an extensive evaluation of several open-source LLMs on evaluation benchmarks for Russian, demonstrating that Vikhr achieves new state-of-the-art results.
2 Related Work
One of the first notable series of generative LLMs for Russian is ruGPT AI Forever (2022); Zmitrovich et al. (2023). The authors created several models trained for the vanilla language modelling task with the sizes of up to 13b. The models were created from the scratch and trained on large Russian corpora. They are able to handle the linguistic nuances of Russian more effectively than multilingual models Muennighoff et al. (2022). Since the training data was mostly in Russian, these models have efficient tokenization, but the lack of multilingual data (e.g. in English) limits their performance. ruGPT models are not instruction tuned.
Gusev (2023) suggests to leverage reasoning capabilities of existing English-oriented LLMs and adapt them to the Russian language by training LoRA adapters. They also create an Alpaca-like set of Russian instruction-output pairs and performed instruction tuning. They have established the Saiga model series, which has a competitive performance and used to be a reasonable choice for off-the-shelf open-source Russian LLM for the past year. However, the tokenizer in theses models is not adapted, so they experience issues with context and computational efficiency.
Tikhomirov and Chernyshev (2023) address these issues in Saiga. In addition to model tuning on Russian data, they also adapt the model tokenizer. They note that improving tokenization helps to both improve the efficiency of the model and its performance while reducing memory consumption. However, during continued pre-training, the authors freeze the model weights except LM heads and token embeddings, which probably results in the suboptimal performance.
In this work, we take advantage of pre-trained English-oriented LLMs, adapt LLM tokenizer for better computational efficiency, leverage continued pre-training on vast Russian-language corpora with regularization for preventing “catastrophic forgetting”, construct a novel extended set of Russian instruction-output pairs, and perform instruction tuning. The created LLM adaptation pipeline along with the data for continued pre-training and instruction tuning enables Vikhr to achieve new state-of-the-art results for Russian, maintain high performance for English, and demonstrate high computational efficiency.
3 LLM Construction Pipeline
The construction of Vikhr starts from one of English-oriented LLMs. In this work, we discuss the Vikhr model based on Mistral 7B. The strong logical and common reasoning capabilities, as well as the extensive world knowledge present in these LLMs provide an excellent starting point for our model. These features partially transfer to Vikhr, enhancing its performance in generating text in Russian. The process of LLM adaptation to Russian starts with the vocabulary adaptation. Then we perform continued pre-training of the LLM on large Russian datasets to mitigate the vocabulary shift and introduce culture specific knowledge. Finally, we perform fine-tuning of Vikhr on a set of instruction-output pairs in Russian.
3.1 Vocabulary Adaptation
Content Length Tokenization Result
Original Sentence 31 Машинное обучение изменяет мир
Mistral Tokenizer 13 [‘Ма’, ‘шин’, ‘ное’, ‘об’, ‘у’, ‘чение’, ‘из’, ‘мен’, ‘я’, ‘ет’, ‘ми’, ‘р’ ]
Vikhr Tokenizer 7 [‘Ма’, ‘шин’, ‘ное’, ‘обучение’, ‘изменяет’, ‘мир’]
Table 1: Tokenizer comparisons between the original Mistral model and Vikhr
Figure 1: The Vikhr tokenizer efficiency in comparison to tokenizers of other models.
The big drawback of English-oriented LLMs is that each Russian word would be split into multiple tokens: a common case is when symbols in the word become an individual tokens (see example in Table 1). This slows down the generation by multiple times, reduces the amount of information that could be stored in the context, and drastically hurts the generation quality.
To mitigate this problem in Vikhr, we adopt the approach suggested in Cui et al. (2023); Tikhomirov and Chernyshev (2023), where authors rebuild the tokenizer using a language-specific corpus. In particular, we trained a SentencePiece tokenizer Kudo and Richardson (2018) with a 40k vocabulary on the RuLM dataset Gusev (2023). As can be seen from Figure 1, the resulting tokenizer for Russian is much more efficient than the tokenizer of the original English-oriented model.
Data Source
Approx. size
Tokens
Scientific papers 20 2.5
News articles 4 1
Wikipedia 25 4
Habr 6 1
Other sources 20 2.5
Table 2: The statistics of the Russian-language datasets for continued pre-training.
3.2 Continued Pre-training
The new vocabulary requires also new embedding matrices and LM heads. The tokens that were present in the original vocabulary are initialized with the old embeddings, the new tokens are initialized by averaging the embeddings of their pieces in the original embedding matrix Hewitt (2021). The similar approach is also applied to LM heads. Training model with these modifications requires much more computational resources than the mainstream technique for adaptation of LLMs to new languages based on LoRA adapters Hu et al. (2021), as it requires to perform continued pre-training of the whole model and on much more language-specific data to mitigate the shift in the vocabulary.
The dataset for continued pre-training is constructed from Russian Wikipedia, news articles, scientific papers, top 100k up-voted posts on Habr, and some other sources. The statistics of these datasets is presented in Table 2. The total number of tokens used for this step is 11 billion.
We note that the continued pre-training of a LLM might partially eliminate the reasoning capabilities present in the original English-oriented model. This drastically affects the model performance. In our preliminary experiments, continued pre-training may result even in worse performance on Russian benchmarks compared to the original model. To alleviate the “catastrophic forgetting”, we use the loss regularization with KL penalty between the probability distribution of Vikhr and the reference English-oriented original LLM:
$L_{\text{Vikhr}}=L_{\text{CE}}+KL\left(P_{\text{Vikhr}}\|P_{\text{Ref }}\right).$ (1)
In practice, we implement this approach using the SLERP interpolation of model losses Goddard et al. (2024).
Hyperparam. Value
LR $1\times 10^{-3}$
AdamW eps $1\times 10^{-8}$
Num warmup steps 10
AdamW betas $0.99$, $0.95$
Accumulation steps $128$
Batch size $3$
Epochs $1$
Sequence length $1024$
Table 3: The hyperparameters for continued pre-training.
To speed up the process of continued pre-training, we use an optimized Flash attention implementation [^2] https://huggingface.co/docs/optimum/bettertransformer/tutorials/convert. As an optimization algorithm, we leverage AdamW as it trades some memory efficiency in favor of robustness to the hyperparameter choice. The hyperparameters used for continued pre-training are presented in Table 3.
3.3 Instruction Tuning
Instruction tuning is an essential step in reaching high zero-shot performance with LLMs. It also allows to obtain more natural communication with the model without complex prompting. Further fine-tuning techniques such as RLHF Ouyang et al. (2022), which require input from the assessors, are also crucial for such tasks as multicriteria alignment. However, the most significant performance gains are still achieved through instruction tuning Jha et al. (2023).
Previously, Gusev (2023) constructed an open-source set of instruction-output pairs for the Russian language (Saiga). The core Saiga dataset was created similar to Alpaca by querying ChatGPT (gpt-3.5-turbo) Taori et al. (2023). In this work, we extend this set by translating two English instruction datasets. First, we translated instructions for the FLAN model Wei et al. (2021) and generated answers in Russian using ChatGPT. Originally, FLAN instructions were constructed automatically from annotated datasets using templates to facilitate multitask and zero-shot capabilities of seq2seq models. Later, it was shown that this data also helps to improve decoder-only chat-oriented models as well. Second, we construct Veles [^3] https://huggingface.co/datasets/Vikhrmodels/Veles-2.5 by translating the English OpenHermes Teknium (2023) instruction dataset. We also include without translation Nectar [^4] https://huggingface.co/datasets/berkeley-nest/Nectar Zhu et al. (2023) – the English instruction dataset. It helps to keep the performance of Vikhr high also for English. Since the majority of the outputs were machine generated there are many low quality outputs. To mitigate this problem, we filtered out low quality pairs using a reward model trained on human data. The statistics of the Vikhr instruction datasets is presented in Table 4.
Instruction Set Language # instances
Veles Russian 30k
Nectar English 50k
Saiga Russian 100k
ruFLAN Russian 500k
Table 4: The statistics of instruction datasets.
Hyperparam. Value
LR $1\times 10^{-5}$
AdamW, eps $1\times 10^{-8}$
Num warmup steps 10
AdamW, betas $0.99$, $0.95$
Accumulation steps $64$
Batch size $3$
Num epochs $3$
Sequence length $1024$
Table 5: The hyperparameters for instruction tuning.
LLM
Pre-train on
Russian
Training
Method
En-MMLU Ru-MMLU CheGeKa
Russian
SuperGLUE
MERA
MTS AI Chat 7B (closed-source) $\diamondsuit$ false sft+dpo - 0.689 0.083 0.56 0.479
GigaChat-7B (closed-source) $\diamondsuit$ true sft+dpo - 0.67 0.451* 0.71* 0.479
aya101 false pt+sft 0.41 0.37 0.005 0.36 0.320
Mistral-7B-Instruct-v0.2 false none 0.60 0.78 0.005 0.57 0.400
rccmsu/ruadapt-mistral-7b-v0.1 false pt+sft 0.61 0.72 0.005 0.64 0.421
rugpt13b true none 0.25 0.25 0.132 0.52 0.208
saiga-mistral-7b-lora false sft 0.60 0.76 0.223 0.64 0.442
saiga-llama3-8b false sft 0.59 0.78 0.225 0.66 0.476
Vikhr-7B-instruct_0.2 true pt+sft 0.62 0.80 0.231 0.67 0.485
Table 6: Evaluation results for Russian and multilingual LLMs. Pre-train on Russian means that the model underwent (continued) pre-training on Russian data. The following abbreviations are used: sft – instruction tuning, pt – (continued) pre-training; dpo – direct preference optimization. $\diamondsuit$ The results for GigaChat and MTS AI are taken from the leaderboards. The best result among open-source models is highlighted with bold, the second best is underscored. The best result among closed-source proprietary models is marked with *.
Contrary to Saiga, we do not use LoRA adapters and just as in the phase of continued pre-training, we update all model parameters. The hyperparameters for the instruction tuning phase are presented in Table 5.
3.4 Hardware
Vikhr was trained on eight NVIDIA A100 GPUs 80GB. We spend approximately 1,000 GPU hours for the continued pre-training phase and 60 hours for instruction tuning.
4 Experiments
4.1 Experimental Setup
Benchmarks.
The evaluation was performed on MMLU Hendrycks et al. (2021), Ru-MMLU [^5] https://github.com/NLP-Core-Team/mmlu_ru, CheGeKa, Russian SuperGLUE Shavrina et al. (2020), and MERA Fenogenova et al. (2024). MMLU (En-MMLU) evaluates LLMs across 57 subjects with multiple-choice questions, assessing a model’s broad knowledge and reasoning abilities. We use this benchmark to verify that the model retains bi-lingual capabilities. In the results, we report the accuracy@1 score. RuMMLU is a translation of MMLU with GPT-3.5 to Russian. Just as for MMLU, we report the accuracy@1 score. CheGeKa is based on questions from the game “What? Where? When?”. This benchmark contains challenging open-ended questions, requiring logical reasoning and world knowledge. It includes 29,376 training and 416 test instances. The reported evaluation metric is the F1 score. Russian SuperGLUE is a benchmark similar to well-known English SuperGLUE Wang et al. (2019). It tests LLMs on various natural language understanding tasks like reading comprehension and textual entailment. The metric reported in the results is accuracy@1. The MERA benchmark encompasses 21 evaluation tasks for generative LLMs in 11 skill domains. Note that among other tasks MERA also includes CheGeKa, RuMMLU, and one of the subtasks of SuperGLUE (RWSD). The reported evaluation metric is the total score, which is the average of scores across all non-diagnostic tasks.
Baselines.
We compare Vikhr to six open-source and two proprietary closed-source competitors of the similar size. Open-source models: aya101 – a massively multilingual LLM from CohereAI that follows instructions in 101 languages [^6] https://huggingface.co/CohereForAI/aya-101, it shows state-of-the-art results among massively multilingual LLMs; Mistral-7B-0.2-instruct – an English-oriented LLM that was used as the base model for Vikhr; rccmsu/ruadapt_mistral_saiga_7b_v0.1 – a Russian-oriented LLM that was constructed from the Mistral model using similar adaptations of the tokenizer, token embeddings, and the LM head Tikhomirov and Chernyshev (2023); saiga-mistral-7b-lora and saiga-llama3-8b – two versions of the Saiga models based on English-oriented LLMs and obtained by fine-tuning LoRA adapters on the Saiga instruction dataset [^7] https://huggingface.co/collections/IlyaGusev. Closed-source proprietary models for Russian: MTS AI Chat [^8] https://huggingface.co/MTSAIR/multi_verse_model and GigaChat-7b. The access to GigaChat weights is closed, so the reported results are taken from the leaderboards [^9] https://mera.a-ai.ru/ru/submits/10257. The results of MTS AI Chat are also taken from the leaderboard [^10] https://mera.a-ai.ru/ru/submits/10290.
4.2 Results
The evaluation results are presented in Table 6. As we can see, Vikhr outperforms all open-source models, including the ones that were built specifically for Russian. It also slightly outperforms its parent model Mistral on the En-MMLU benchmark, which might be the result of longer pre-training. The second place with close scores for all 4 Russian language benchmarks is obtained by the Saiga model based on recently released Llama-3. The high scores of this model probably are the result of the transfer of the outstanding performance of Llama-3. Since Saiga based on Llama-3 outperforms Saiga based on Mistral, we expect that applying our adaptation pipeline to Llama-3 would also help further improving the state of the art.
We note that the original Mistral-7B-0.2-instruct, despite being an English-oriented model, demonstrates competitive performance in 3 out of 4 Russian benchmarks. This indicates demonstrates that such models could be viable alternatives. The only dataset, where its performance is very low is CheGeKa, which is related to open-ended question-answering. This may be due to the lack of culture-specific knowledge, as the English-oriented model has not seen much Russian texts. Note that the MTS AI Chat also shows very low results on CheGeKa, which might also indicate the lack of culture-specific knowledge.
The proprietary model GigaChat substantially outperforms Vikhr on CheGeKa and notably on Russian SuperGLUE. We assume this is due to the use of much larger Russian datasets for pre-training. However, surprisingly, it falls behind Vikhr on Ru-MMLU. On all benchmarks, Vikhr outperforms the the proprietary competitor from MTS AI.
5 Conclusion
We have presented Vikhr – a new state-of-the-art open-source instruction-following LLM oriented on the Russian language. To create Vikhr, we developed a comprehensive pipeline for adapting English-oriented LLMs to Russian. The pipeline includes the adaptation of the tokenizer vocabulary, continued pre-training of the entire model, and instruction tuning. We have also constructed a new dataset for instruction tuning by expanding the Saiga dataset with automatically translated and cleaned English instruction datasets. Our extensive work enabled Vikhr to outperform the known baselines, while maintaining computational efficiency.
We hope that the published models will foster the research on LLMs and enhance the diversity of languages incorporated into research agendas.
Limitations
We do not introduce additional restrictions to the usage of our models. However, the users must comply with the license of the base model and instruction datasets.
We do not implement RLHF / DPO fine-tuning of Vikhr due to the lack of the resources for human annotation. We expect further performance improvements from these techniques.
We do not introduce additional instruction-output pairs to facilitate LLM alignment. However, we note that the majority of the data for supervised fine-tuning of Vikhr are obtained from the ChatGPT model series, so our model partially inherits its alignment.